OpenStudy (anonymous):

differentiate 3^(1/2)Sec^(-1)(2xy)

OpenStudy (anonymous):

hi maleck use the formula Dx(sec^-1 u)= +-(1/(usqrt(u^2 -1))) , let u=2xy, therefore differentiate 3^(1/2)Sec^(-1)(2xy) =3^(1/2) [+-(2(xdy/dx +y))/(2xysqrt(4x^2 y^2 -1))

OpenStudy (anonymous):

hi maleck use the formula Dx(sec^-1 u)= +-(1/(usqrt(u^2 -1))) , let u=2xy, therefore differentiate 3^(1/2)Sec^(-1)(2xy) =3^(1/2) [+-(2(xdy/dx +y))/(2xysqrt(4x^2 y^2 -1))